ohig?? - b17172

Katherine: if enough space, continue recording onto the 8TB module (as is AOV). The 2 TB module is for the R4. (Lucia)


Hobart 12m:

Disk VSN: HOB+0125/32000/1024

Data volume at beginning: 8170.507 GB

  • Experiment started okay (Patrick)

Katherine 12m:

Disk VSN: GSFC+011/8000

Data volume at beginning: 0.0 GB

  • 1827UT: Experiment started late due to network being down. (Patrick)
  • Module has no VSN assigned, however the label on the module is correct. Need to set module VSN. (Patrick)
  • Weather command 'wx' is not working, only returns 19.09, 998.8990. 56.20. (Patrick)
  • 00:30UT - Problems when inserting a new module. Mk5 locked up. Jamie fixed it. Mk5 time was off, had to resync (corrected for 1 second). Lost scans 0011 through 0037. (Lucia)
  • 00:40UT - Weather is now working again (Lucia).
  • 06:50UT - weather sensor is stuck again, gives identical value each scan. Jamie is fixing it, though he still has to write down how. (Lucia)

Yarragadee 12m:

Disk VSN: HART+104/8000/1024

Data volume at beginning: 37.949 GB

  • 1755UT: Experiment started late because antenna was stuck. e-RemoteCtrl won't connect, streaming logs instead (Patrick)
  • 1859UT: Successfully started e-RemoteCtrl. Left log streaming (Patrick)